# When Measurement Mediates the Effect of Interest

### v2

@@ 3.3 Two-Stage TMLE for Estimation and Inference
Two-Stage TMLE is a general approach to identifying and estimating causal effects in a settings with missing data and dependent data [27, 28]. In the first stage, we fully stratify on each cluster and focus on controlling for differential missingness/censoring of individual-level data. In the second stage, we use estimates from the first stage to evaluate the total effect and focus on improving efficiency. In the following, we propose an implementation of Two-Stage TMLE to address unique challenges when evaluating the total effect of strategies to improve reach into the target population and outcomes among those reached. As detailed below, fully stratifying on each cluster in the first stage simplifies the multilevel-mediation-missingness problem (Figure 1). Therefore, we can control for differential measurement without losing the effect of the cluster-level exposure (i.e., without blocking the indirect effect).
@@ 3.4 Stage 1: Identifying and estimating in each cluster
In Stage 1, we consider each cluster in turn. Within each cluster, the cluster-level covariates and intervention are constant. Therefore, the observed data for cluster $j$ simplify to $O_{ij}=(W_{ij},\Delta_{ij},Y1_{ij},Y2_{ij})$ for $i=1,\ldots,N_{j}$. Equally importantly, our Stage 1 causal estimand simplifies to the counterfactual probability of the outcome for persons in the target population: $Y^{c*}$= $\mathbb{P}(Y2=1\mid Y1^{*}=1)$. In OPAL, this is the probability of initiating PrEP among persons with HIV risk. Crucially, stratifying on cluster allows the missingness mechanism to vary by cluster, while capturing the impacts of the cluster-level variables, including the intervention $A^{c}$.
Following [45, 26, 27, 28], we re-express this causal estimand as the joint probability of being in the target population and having the outcome, divided by the probability of being in the target population:
@@
In OPAL, $Y^{c*}$ is now written as the joint probability of initiating PrEP and being at risk of HIV acquisition, divided by the prevalence of HIV risk. The numerator and denominator can then be identified and estimated, separately.
We first focus on the numerator. Throughout, having the outcome is contingent on being in the target population (e.g., starting PrEP is contingent upon on testing HIV-negative and reporting risk). Therefore, the numerator simplifies to the observed proportion with the outcome: $\mathbb{P}(Y2=1)$. This framework naturally extends to more complex settings where the outcome is also subject to missingness [26, 27, 28].
We now turn to the denominator. Throughout, the denominator defines the target population and is subject to missingness: $\mathbb{P}(Y1^{*}=1)$. To identify the denominator from the observed data distribution, we need to consider the plausibility of various missing data assumptions. Suppose the missing-completely-at-random (MCAR) held within each cluster [3]. For a given cluster in OPAL, the corresponding assumption would be that HIV risk among persons who go to the clinic for screening is perfectly representative of HIV risk among persons who do not. This is highly implausible, and we, instead, consider the missing-at-random (MAR) assumption [3]: for each cluster, HIV risk among those screened is representative of risk among those missed within all possible values of the baseline covariates $W$. Importantly, the adjustment set for a given individual can include covariates of other cluster members (e.g., their friends). Furthermore, in OPAL where a cluster is a group of nearby drinking venues, the adjustment set can also include characteristics of the venue (e.g., type of alcohol served or presence of on-site rooms for sex work). This missing data assumption can be represented as $Y1^{*}\perp\Delta\mid W$ and is further relaxed in Section 5 to adjust for time-varying covariates, which are potentially impacted by the exposure $A^{c}$. Additionally, we require that the probability of being measured within covariate strata be bounded away from 0: $\mathbb{P}(\Delta=1\mid W=w)>0$ for all possible $w$. Under these assumptions, the statistical estimand for the denominator is given by the G-computation formula [54]: $\mathbb{E}[\mathbb{E}(Y1\mid\Delta=1,W)]$.
@@
In OPAL, $Y^{c}$ is interpreted as the probability of starting PrEP, divided the adjusted prevalence of HIV risk. To estimate the numerator, we use the empirical proportion with the outcome. To estimate the denominator, we use TMLE, a doubly robust, substitution estimator optimizing the bias-variance trade-off for the target parameter [55]. Within TMLE, we apply the Super Learner, an ensemble machine learning algorithm [56], for flexible estimation of the the outcome regression $\mathbb{E}(Y1\mid\Delta=1,W)$ and measurement mechanism $\mathbb{P}(\Delta=1\mid W)$. Briefly, the TMLE algorithm updates the initial estimates of $\hat{\mathbb{E}}(Y1\mid\Delta=1,W)$ by using information from ^$\hat{\mathbb{P}}(\Delta=1\mid W)$ and obtains targeted estimates $\hat{\mathbb{E}}^{\star}(Y1\mid\Delta=1,W)$, which are averaged across individuals.
Since we fully stratify on cluster in Stage 1, we need to implement the estimation algorithm $J$ times to obtain estimates for each cluster $j=\{1,\ldots,J\}$:
@@
To obtain confidence intervals for these estimates $\hat{Y}^{c}$, we apply the Delta method to the influence curve of the numerator and denominator.
@@ 3.5 Stage 2: Identifying and estimating the total effect
Stage 1 focuses on identifying and estimating the expected outcome among the underlying target population for each cluster. However, recall our primary research question is the total effect of the health promotion strategy: $\Psi^{*}=\mathbb{E}[Y^{c*}(1)-Y^{c*}(0)]$. In OPAL, $\Psi^{*}$ is the expected difference in the counterfactual uptake of PrEP (among persons with HIV risk) if all clusters received multi-disease messaging $Y^{c}$(1) versus HIV-focused messaging $Y^{c}$(0). In Stage 2, we focus on evaluating the total effect $\Psi^{*}$ with maximum precision. In Stage 2, the observed data are at the cluster-level and given by $O^{c}=(E^{c},W^{c},A^{c},\hat{Y}^{c})$, where $W^{c}$ denote aggregates of the individual-level baseline covariates W and $\hat{Y}^{c}$ are the cluster-level endpoint estimates from Stage 1.
@@
$\hat{\Psi}_{unadj}=\frac{1}{J}\sum_{j=1}^{J}\mathbbm{1}(A^{c}=1)(\hat{Y}^{c}_{j})-\mathbbm{1}(A^{c}=0)(\hat{Y}^{c}_{j})$ (1)
@@
However, statistical precision and power can be improved by adjusting for baseline covariates that are predictive of the outcome [57, 58, 59, 60, 61, 44, 27]. Therefore, we implement a cluster-level TMLE to obtain a more efficient estimate of the total effect. Briefly, an initial estimate of the cluster-level outcome regression $\hat{\mathbb{E}}(\hat{Y}^{c}\mid A^{c}=a^{c},E^{c},W^{c})$ is updated using information in the cluster-level propensity score ^$\hat{\mathbb{P}}(A^{c}=1\mid E^{c},W^{c})$. Then targeted estimates of the expected outcome under each trial arm are averaged to obtain a point estimate:
$\hat{\Psi}_{tmle}=\frac{1}{J}\sum_{j=1}^{J}\hat{\mathbb{E}}^{\star}(\hat{Y}^{c}\mid A^{c}=1,E^{c}_{j},W^{c}_{j})-\hat{\mathbb{E}}^{\star}(\hat{Y}^{c}\mid A^{c}=0,E^{c}_{j},W^{c}_{j})$ (2)
@@
In a trial setting, we implement TMLE with Adaptive Pre-specification to data-adaptively select the adjustment approach that maximizes empirical efficiency [62, 63]. In this procedure, we pre-specify candidate algorithms for estimation of the outcome regression and the known propensity score. In trials with few clusters, we recommend limiting these algorithms to “working” generalized linear models (GLMs) adjusting for a single covariate. In trials with many clusters, we can expand the candidate set to include more adaptive approaches adjusting for multiple covariates (e.g., stepwise regression and multivariate adaptive regression splines). In addition, we pre-specify the loss function to be the squared of the influence curve of the cluster-level TMLE; the expectation of this loss function is the variance of the TMLE. Finally, we pre-specify a cross-validation scheme to evaluate performance and collaboratively select the combination of outcome regression and propensity score estimators that minimize the cross-validated risk estimate. (The algorithm is “collaborative” because the propensity score is fit in response to estimation of the outcome regression [64]) In the event that there are no gains in efficiency with any of the candidate adjustment approaches, this procedure defaults to the unadjusted estimator.
To obtain statistical inference, we obtain a variance estimate by taking the sample variance of the estimated influence curve divided by the number of independent units $J$. To construct Wald-type $95\%$ confidence intervals and test the null hypothesis, we use the Student’s t-distribution as small sample approximation to the Gaussian distribution [65]. This inferential approach relies on Two-Stage TMLE being an asymptotically linear estimator [66]. Specifically, under the conditions detailed in Balzer et al. [27], we have that the cluster-level TMLE minus the true statistical estimand behave in first order as an empirical mean of a mean zero and finite variance function known as the influence curve i.e. $\hat{\Psi}(a^{c})-\Psi(a^{c})=\frac{1}{J}\sum^{J}_{j}IC_{j}+R_{J}$ where $IC_{j}$ is the cluster-level influence curve and $R_{J}$ is a second order remainder term that goes to zero in probability. Briefly, we need that Stage 1 estimation of $Y^{c}$ has a negligible contribution to the remainder term $R_{J}$ and that Stage 2 estimators for the outcome regression and propensity score satisfy the usual regularity conditions [60]. These Stage 2 conditions are naturally satisfied when using Adaptive Pre-specification to select among working GLMs.
@@
Altogether, Two-Stage TMLE is a multiply robust estimator. In Stage 1, TMLE offers a doubly robust estimation of $Y^{c}$ in each of the $J$ clusters; if either outcome regression or measurement mechanism is consistently estimated, we will have a consistent estimate of the cluster-level endpoint $Y^{c}$. Then in Stage 2, TMLE offers a model-robust procedure to maximize efficiency when evaluating the total effect of the cluster-level strategy. Importantly, bias and variability in the Stage 1 estimates $\hat{Y}^{c}$ contribute directly to the cluster-level influence curve and, thus, variance in Stage 2.
@@ 4 Simulation Study
To examine the finite sample performance of our proposed approach, we conducted a simulation study, reflecting the multilevel-mediation-missing data problem. In these simulations, we made measurement (the mediator) strongly dependent on the baseline covariates and the randomized intervention, while setting the direct effect of the intervention on the outcome to be very weak. To reflect OPAL, we focused on a trial with $J=150$ clusters. Simulations were conducted in R version 4.4.1 Code to replicate these simulations is available at
@@ 4.1 Generating the Data
For each cluster $j=\{1,\ldots,J\}$, we repeated the following data generating process. First, we generated cluster-level latent variables $U_{E1^{c}}$ $\sim$ Uniform(0,1) and $U_{E2^{c}}$ $\sim$ Uniform(0,0.5) and then the cluster-level observed covariates $E1^{c}$ $\sim$ Normal($U_{E1^{c}}$,1) and $E2^{c}$ $\sim$ Normal($U_{E2^{c}}$,1). Next, we generated the number of individuals in the cluster as $N\sim$Normal(100,10), rounding to the nearest whole number. For the $N$ individuals in the cluster, we independently generated baseline covariates: $W1$ $\sim$ Uniform(18,60) (scaled), $W2\sim$ Binomial(0.6), and $W3\sim$ Binomial(0.65). We aggregated these to the cluster-level by taking the empirical mean to obtain $W1^{c}$, $W2^{c}$, and $W3^{c}$, respectively. Then the cluster-level intervention $A^{c}$ was randomly assigned: $A^{c}=\mathbf{1}(U_{A^{c}}<0.5)$ where $U_{A^{c}}\sim$ Uniform(0,1).
We generated the underlying indicator of being in the target population $Y1^{*}$ for each individual as a simple function of the covariates but not the intervention: $Y1^{*}=\mathbf{1}[U_{Y1^{*}}<logit^{-1}(-0.8+0.3E1^{c}+0.15E2^{c}+2W1+1.5W2-3.0W3+0.55U_{E1^{c}}+0.15U_{E2^{c}})]$ with $U_{Y1^{*}}\sim$ Uniform(0,1). We generated the corresponding measurement indicator as a complex function of the covariates and the randomized intervention: $\Delta=\mathbf{1}(U_{\Delta}<p_{\Delta})$ where $U_{\Delta}\sim$ Uniform(0,1) and
$p_{\Delta}=W3\times logit^{-1}$ $(0.1+0.8A^{c}-2W1+0.8W2\times A^{c}+0.75W3$
$-0.75E1^{c}+0.75E2^{c}-0.15U_{E1^{c}}+0.15U_{E2^{c}})$
$+(1-W3)\times$ $logit^{-1}(0.1+0.8A^{c}+2W1-0.8W2\times A^{c}-0.75W3$
$+0.75E1^{c}-0.75E2^{c}+0.15U_{E1^{c}}-0.15_{UE2^{c}})$
Therefore, the impact of the covariates on measurement (i.e., health screening) was highly dependent on $W3$. Additionally, since measurement $\Delta$ was a function of the cluster-level exposure $A^{c}$, we also generated counterfactual mediator values $\Delta(a^{c})$ by setting $A^{c}=\{0,1\}$ for all individuals.
Finally, we generated the individual-level outcome $Y2$ as a function of the baseline covariates, randomized intervention, and measurement/screening status. We set the outcome deterministically to zero for individuals who were not screened or screened but not part of the target population: $Y2=0$ if $\Delta=0$ or if $Y1=\Delta Y1^{*}=0$. Among those screened and in the target population ($Y1=\Delta Y1^{*}=1$), we generated the outcome as $Y2=\mathbf{1}[U_{Y2}<logit^{-1}(0.2+0.08A^{c}+0.5W1+0.4W2-2E1^{c}+0.87E2^{c}-0.08A^{c}W_{3})]$ with $U_{Y2}\sim$ Uniform(0,1). Thus, the direct effect of the randomized exposure $A^{c}$ on the outcome $Y2$ was weak. We also generated the counterfactual outcomes $Y2(a^{c})$ setting $A^{c}=\{0,1\}$, but not setting $\Delta=1$. Instead, through the above process, the counterfactual outcome $Y2(a^{c})$ was a function of the counterfactual mediator $\Delta(a^{c})$. We calculated the cluster-level counterfactual outcome $Y^{c*}(a^{c})=\mathbb{P}[Y2(a^{c})=1\mid Y1^{*}=1]$
and evaluated the total effect $\Psi^{*}=\mathbb{E}[Y^{c*}(1)-Y^{c*}(0)]$ with difference in average counterfactual cluster-level outcomes for a population of 5000 clusters.

### v3

@@ 3.3 Two-Stage TMLE for Estimation and Inference
Two-Stage TMLE is a general approach to identifying and estimating causal effects in a settings with missing data and dependent data [4, 38]. In the first stage, we stratify on each cluster and focus on controlling for differential missingness/censoring of individual-level data. In the second stage, we use estimates from the first stage to evaluate the total effect and focus on improving efficiency. In the following, we propose an implementation of Two-Stage TMLE to address unique challenges when evaluating the total effect of strategies to expand health screening and to improve health outcomes. As detailed below, fully stratifying on each cluster in the first stage simplifies the multilevel-mediation-missingness problem (Figure 1). Specifically, we can control for differential screening without losing the effect of the cluster-level exposure (i.e., without blocking the indirect effect).
@@ 3.4 Stage 1: Identifying and estimating in each cluster
In Stage 1, we consider each cluster in turn. Within each cluster, the cluster-level covariates and intervention are constant. Therefore, the observed data for cluster $j$ simplify to $O_{ij}=(W_{ij},\Delta_{ij},Y1_{ij},Y2_{ij})$ for $i=1,\ldots,N_{j}$. Equally importantly, our Stage 1 causal estimand simplifies to the counterfactual probability of the outcome for persons in the focus population: $Y^{c*}$= $\mathbb{P}(Y2=1\mid Y1^{*}=1)$. In OPAL, this is the probability of initiating PrEP among persons with HIV risk. Crucially, stratifying on cluster allows the missingness mechanism to vary by cluster, while capturing the impact of the cluster-level variables, including the intervention $A^{c}$.
Following [9, 6, 4, 38], we re-express this causal estimand as the joint probability of being in the focus population and having the outcome, divided by the probability of being in the focus population:
@@
In OPAL, $Y^{c*}$ is now written as the joint probability of initiating PrEP and being at risk of HIV acquisition, divided by the prevalence of HIV risk.
The numerator and denominator can then be identified and estimated, separately. We first focus on the numerator. Throughout, having the outcome is contingent on being in the focus population. In OPAL, starting PrEP is contingent upon on testing HIV-negative and reporting risk. Therefore, the numerator simplifies to the observed proportion with the outcome: $\mathbb{P}(Y2=1)$.
This framework naturally extends to more complex settings where the outcome is also subject to missingness [6, 4, 38].
We now turn to the denominator. Throughout, the denominator defines the focus population and is subject to missingness: $\mathbb{P}(Y1^{*}=1)$. To identify the denominator from the observed data distribution, we need to consider the plausibility of various missing data assumptions. Suppose the missing-completely-at-random (MCAR) held within each cluster [50]. For a given cluster in OPAL, the corresponding assumption would be that HIV risk among persons who go to the clinic for screening is perfectly representative of HIV risk among persons who do not. This is highly implausible, and we, instead, consider the missing-at-random (MAR) assumption [50]: for each cluster, HIV risk among those screened is representative of risk among those missed within all possible values of the baseline covariates $W$. Importantly, the adjustment set for a given individual can include covariates of other cluster members (e.g., their friends). Furthermore, the adjustment set for a given individual can include covariates of the sub-cluster (e.g., their household). In OPAL where a cluster is a group of nearby alcohol-serving venues, the adjustment set will include characteristics of the venue (e.g., type of alcohol served or presence of on-site rooms for sex work). This missing data assumption can be represented as $Y1^{*}\perp\Delta\mid W$ and is further relaxed in Section 5 to adjust for time-varying covariates, which are potentially impacted by the exposure $A^{c}$. Additionally, we require that the probability of being measured within covariate values be bounded away from 0: $\mathbb{P}(\Delta=1\mid W=w)>0$ for all possible $w$. Under these assumptions, the statistical estimand for the denominator is given by the G-computation formula [47]: $\mathbb{E}[\mathbb{E}(Y1\mid\Delta=1,W)]$.
@@
In OPAL, $Y^{c}$ is interpreted as the probability of starting PrEP, divided the adjusted prevalence of HIV risk. To estimate the numerator, we use the empirical proportion with the outcome. To estimate the denominator, we use TMLE, a doubly robust, substitution estimator optimizing the bias-variance trade-off for the target parameter [66]. Within TMLE, we apply the Super Learner, an ensemble machine learning algorithm [65], for flexible estimation of the the outcome regression $\mathbb{E}(Y1\mid\Delta=1,W)$ and measurement mechanism $\mathbb{P}(\Delta=1\mid W)$. Briefly, the TMLE algorithm updates the initial estimates of $\hat{\mathbb{E}}(Y1\mid\Delta=1,W)$ by using information from $\hat{\mathbb{P}}(\Delta=1\mid W)$ and obtains targeted estimates $\hat{\mathbb{E}}^{\star}(Y1\mid\Delta=1,W)$, which are averaged across individuals. Since we stratify on each cluster in Stage 1, we need to implement the estimation algorithm $J$ times to obtain estimates for each cluster $j=\{1,\ldots,J\}$:
@@ 3.5 Stage 2: Identifying and estimating the total effect
Stage 1 focuses on identifying and estimating the expected outcome among the underlying focus population for each cluster. However, recall our primary research question is the total effect of the health strategy: $\Psi^{*}=\mathbb{E}[Y^{c*}(1)-Y^{c*}(0)]$. In OPAL, $\Psi^{*}$ is the expected difference in the counterfactual uptake of PrEP (among persons with HIV risk) if all clusters received multi-disease messaging $Y^{c}$(1) versus HIV-focused messaging $Y^{c}$(0). In Stage 2, we focus on evaluating the total effect $\Psi^{*}$ with maximum precision. In Stage 2, the observed data are at the cluster-level and given by $O^{c}=(E^{c},W^{c},A^{c},\hat{Y}^{c})$, where $W^{c}$ denote aggregates of the individual-level baseline covariates W and $\hat{Y}^{c}$ are the cluster-level endpoint estimates from Stage 1.
@@
$\hat{\Psi}_{unadj}\ =\ \hat{\mathbb{E}}(\hat{Y}^{c}|A=1)-\hat{\mathbb{E}}(\hat{Y}^{c}|A=0)\ =\ \frac{\sum_{j=1}^{J}\mathbbm{1}(A^{c}_{j}=1)\hat{Y}^{c}_{j}}{\sum_{j=1}^{J}\mathbbm{1}(A^{c}_{j}=1)}-\frac{\sum_{j=1}^{J}\mathbbm{1}(A^{c}_{j}=0)\hat{Y}^{c}_{j}}{\sum_{j=1}^{J}\mathbbm{1}(A^{c}_{j}=0)}$ (1)
@@
However, statistical precision and power can be improved by adjusting for baseline covariates that are predictive of the outcome [17, 58, 15, 36, 48, 11, 4]. Therefore, we implement a cluster-level TMLE to obtain a more efficient estimate of the total effect. Briefly, an initial estimate of the cluster-level outcome regression $\hat{\mathbb{E}}(\hat{Y}^{c}\mid A^{c}=a^{c},E^{c},W^{c})$ is updated using information in the cluster-level propensity score $\hat{\mathbb{P}}(A^{c}=1\mid E^{c},W^{c})$. Then targeted estimates of the expected outcome under each trial arm are averaged to obtain a point estimate:
$\hat{\Psi}_{tmle}=\frac{1}{J}\sum_{j=1}^{J}\hat{\mathbb{E}}^{\star}(\hat{Y}^{c}\mid A^{c}=1,E^{c}_{j},W^{c}_{j})-\frac{1}{J}\sum_{j=1}^{J}\hat{\mathbb{E}}^{\star}(\hat{Y}^{c}\mid A^{c}=0,E^{c}_{j},W^{c}_{j})$ (2)
@@
In a trial setting, we implement TMLE with Adaptive Pre-specification to data-adaptively select the adjustment approach that maximizes empirical efficiency [7, 3, 8]. In this procedure, we pre-specify candidate algorithms for estimation of the outcome regression and the known propensity score. In trials with few clusters, we recommend limiting these algorithms to “working” generalized linear models (GLMs) adjusting for a single covariate [48]. In trials with many clusters, we can expand the candidate set to include more adaptive approaches adjusting for multiple covariates (e.g., stepwise regression and multivariate adaptive regression splines). The unadjusted estimator is always included as candidate. In addition, we pre-specify the loss function to be the square of the influence curve of the cluster-level TMLE; the expectation of this loss function is the variance of the TMLE [49]. Finally, we pre-specify a cross-validation scheme to evaluate performance and collaboratively select the combination of outcome regression and propensity score estimators that minimize the cross-validated risk estimate. (The algorithm is “collaborative” because the propensity score is fit in response to estimation of the outcome regression [57].) In the event that there are no gains in efficiency with any of the candidate adjustment approaches, this procedure defaults to the unadjusted estimator.
3.6 Statistical Inference
@@
Two-Stage TMLE is a multiply robust estimator. In Stage 1, TMLE offers a doubly robust estimation of $Y^{c}$ in each of the $J$ clusters; if either outcome regression or measurement mechanism is consistently estimated, we will have a consistent estimate of the cluster-level endpoint $Y^{c}$. Then in Stage 2, TMLE offers a model-robust procedure to maximize efficiency when evaluating the total effect of the cluster-level strategy. Importantly, bias and variability in the Stage 1 $\hat{Y}^{c}$ estimates contribute directly to the cluster-level influence curve and, thus, variance in Stage 2.
Under conditions, Two-Stage TMLE is an asymptotically linear estimator [64] — the cluster-level TMLE minus the statistical estimand behaves in first order as an empirical mean of a mean-zero and finite variance function known as the influence curve: $\hat{\Psi}-\Psi=\frac{1}{J}\sum^{J}_{j}IC_{j}+R_{J}$. Here, $IC_{j}$ is the influence curve for cluster $j$ and $R_{J}$ is a second order remainder term that goes to zero in probability. Then we obtain a variance estimate by taking the sample variance of the estimated influence curve divided by the number of independent units $J$. To construct Wald-type $95\%$ confidence intervals and test the null hypothesis, we use the Student’s t-distribution with J2J-2 degrees of freedom as small sample approximation to the Gaussian distribution [23].
The conditions for Two-Stage TMLE to be asymptotically linear are detailed in Balzer et al. [4]. First, we require that Stage 1 estimation of $Y^{c}$ has a negligible contribution to the remainder term $R_{J}$. This holds under the following conditions: the within cluster dependence is weak enough such that the Central Limit Theorem applies in $N_{j}$; the ratio of the number of independent units to the smallest cluster size goes to zero (i.e., $J/min_{j}(N_{j})\rightarrow 0$), and the individual-level estimators of the outcome regression and propensity score converge to their targets at fast enough rates such that the corresponding Stage 1 remainder term is $o_{p}(N_{j}^{-1/2})$. Second, we require that Stage 2 estimators for the cluster-level outcome regression and measurement mechanism satisfy the usual regularity conditions [4]. These Stage 2 conditions are naturally satisfied when using Adaptive Pre-specification to select among working GLMs [36]. While these conditions are asymptotic, our simulation study, described next, evaluates finite sample performance.
@@ 4 Simulation Study
To examine the finite sample performance of our proposed approach, we conducted a simulation study reflecting the multilevel-mediation-missing data problem. In line with a systematic review revealing that most CRTs simply exclude participants with missing data [14], we considered approaches that restricted to participants who screened ($\Delta=1$). In line with common practice when evaluating HIV treatment/prevention strategies [62, 63, 61], we also considered several approaches that simply restricted to persons known to be in the focus population ($Y1=1$). For comparison to more traditional methods, we also considered two “Single-Stage” approaches that simultaneously account for missing data and estimate the intervention effect — pooling individuals across clusters: generalized estimating equations (GEE) and the standard TMLE for clustered data [53, 30]. The considered estimators are described below in detail. Simulations were conducted in R version 4.4.1. Code to replicate these simulations is available at https://github.com/jznakato/multi-level-mediation-missingness
@@ 4.1 Generating the Data
For each cluster $j=\{1,\ldots,J\}$, we repeated the following data generating process. First, we generated cluster-level latent variables $U_{E1^{c}}$ and $U_{E2^{c}}$ by drawing independently from a Unif(1,1)Unif(-1,1) and then the cluster-level observed covariates as $E1^{c}\sim Norm(U_{E1^{c}},.5)$ and $E2^{c}\sim Norm(U_{E2^{c}},.5)$. Next, we generated the number of individuals in the cluster as $N\sim Norm(200,10)$, rounding to the nearest whole number. For the $N$ individuals in the cluster, we independently generated baseline covariates: $W1\sim Norm(0,.5)$, $W2\sim Norm(0,.5)$, and $W3\sim Bin(.5)$. We generated the underlying indicator of being in the focus population $Y1^{*}$ for each individual as $Y1^{*}=\mathbf{1}[U_{Y1^{*}}<logit^{-1}(.5+W1+W2-W3+0.25E1^{c}+0.25E2^{c})]$ with $U_{Y1^{*}}\sim Unif(0,1)$. For each individual, we generated the counterfactual measurement indicator under $a^{c}=\{0,1\}$ as $\Delta(a^{c})=\mathbf{1}[U_{\Delta}<logit^{-1}(-.5+.6a^{c}+.5W1+.4W2-.4a^{c}W3+.4(1-a^{c})W3+.1U_{E1^{c}}+.1U_{E2^{c}})]$ with $U_{\Delta}\sim Unif(0,1)$. Likewise, for each individual, we generated the counterfactual outcome $Y2(a^{c})$ under $a^{c}\in\{0,1\}$ as follows. We set the outcome deterministically to zero if not in the focus population, $Y1^{*}=0$, or if not screened, $\Delta(a^{c})=0$. Otherwise, we generated the counterfactual outcome as $Y2(a^{c})=\mathbf{1}[U_{Y2}<logit^{-1}(.2+.1A^{c}+W1+.5W2+2W3+.2E1^{c}+.2E2^{c})]$ with $U_{Y2}\sim Unif(0,1)$.
Through this process, we generated each cluster’s “full data”, consisting of the baseline covariates $(E1^{c},E2^{c},\bm{W1},\bm{W2},\bm{W3})$, the indicators of being in the focus population $\bm{Y1}^{*}$, and the counterfactual indicators of screening and having the outcome $(\bm{\Delta}(a^{c}),\bm{Y2}(a^{c}))$ for $a^{c}\in\{0,1\}$. We converted the full data to the observed data as follows. We randomized the trial arm $A^{c}$ such that $J/2$ clusters were intervention ($A^{c}=1$) and $J/2$ were control $(A^{c}=0)$. We then assigned the observed indicator for measurement as $\bm{\Delta}=\bm{\Delta}(a^{c})$ when $A^{c}=a^{c}$, the observed indicator for being in the focus population $\bm{Y1}=\bm{Y1}^{*}\times\bm{\Delta}$, and the observed indicator for having the outcome as $\bm{Y2}=\bm{Y2}(a^{c})$ when $A^{c}=a^{c}$.
We repeated this process for 1000 simulated trials each of $J=\{20,30,50,70\}$ clusters. We calculated the true value of the total effect for a population of 5000 clusters as $\Psi^{*}=\mathbb{E}[Y^{c*}(1)-Y^{c*}(0)]$ with the cluster-level counterfactual outcome $Y^{c*}(a^{c})=\mathbb{P}[Y2(a^{c})=1\mid Y1^{*}=1]$. We also simulated under the null by setting the coefficients for $a^{c}$ to be zero in the above data generating process.